# Energy-Efficient Flying LoRa Gateways: A Multi-Agent Reinforcement Learning Approach

### v1

@@ Abstract
With the rapid development of next-generation Internet of Things (NG-IoT) networks, the increasing number of connected devices has led to a surge in power consumption. This rise in energy demand poses significant challenges to resource availability and raises sustainability concerns for large-scale IoT deployments. Efficient energy utilization in communication networks, particularly for power-constrained IoT devices, has thus become a critical area of research. In this paper, we deployed flying LoRa gateways (GWs) mounted on unmanned aerial vehicles (UAVs) to collect data from LoRa end devices (EDs) and transmit it to a central server. Our primary objective is to maximize the global system energy efficiency (EE) of wireless LoRa networks by joint optimization of transmission power (TP), spreading factor (SF), bandwidth (W), and ED association. To solve this challenging problem, we model the problem as a partially observable Markov decision process (POMDP), where each flying LoRa GW acts as a learning agent using a cooperative Multi-Agent Reinforcement Learning (MARL) approach under centralized training and decentralized execution (CTDE). Simulation results demonstrate that our proposed method, based on the multi-agent proximal policy optimization (MAPPO) algorithm, significantly improves the global system EE and surpasses the conventional MARL schemes.
@@ I Introduction
The next-generation Internet of Things (NG-IoT) technologies for fifth-generation (5G) and sixth-generation (6G) applications are revolutionizing communication by enabling seamless data exchange between devices and networks. This evolution is driving intelligent applications in areas such as healthcare, smart cities, agriculture, and autonomous vehicles [1]. With IoT device connections expected to reach 125 billion by 2030 [2], energy consumption poses a significant challenge, particularly in maintaining low-power and long-range communication networks. Enhancing energy efficiency (EE) is therefore essential to align with global sustainability objectives, including the United Nations Sustainable Development Goal 7: Affordable and Clean Energy [3].
Consequently, low-power wide area networks (LPWANs), particularly long-range (LoRa) technology, have emerged as a key solution to enable energy-efficient communication over long distances. However, because of the extensive interconnectivity, these networks are becoming denser, thus increasing the need for efficient resource allocation techniques. Recent studies on resource allocation approaches, especially in cellular communications, have focused on the use of unmanned aerial vehicles (UAVs) to enhance network coverage and efficiency. Although these approaches have been beneficial in traditional cellular networks, their application to IoT-specific systems has been limited. This limitation has led to inefficiencies, particularly as IoT networks face growing traffic and dynamic communication demands. Addressing these challenges requires the development of efficient resource management algorithms that optimize EE while meeting the dynamic requirements of IoT networks.
A common approach in the literature is alternating optimization (AO), which involves formulating an optimization problem, breaking it down into manageable subproblems, and subsequently developing algorithmic solutions to solve each part [4, 5]. However, this approach becomes particularly critical when considering the complexities introduced by the dense interconnectivity and varying traffic demands of IoT-enabled systems such as flying LoRa gateways (GWs). Moreover, a substantial number of recent studies have used machine learning (ML) techniques to enhance EE in LoRa networks. For example, the authors in [6] address the challenges of LPWAN by using Q-learning with multiple agents framework to dynamically allocate transmission power (TP) and spreading factor (SF) in a LoRa network, thus improving EE. In [7], Q-learning is applied to enhance the packet delivery ratio (PDR) by rewarding successfully received packets. However, reliance on Q-tables makes Q-learning impractical for complex optimization tasks. To address this, proximal policy optimization (PPO) was introduced as a more stable and data-efficient alternative [8]. In [9], the authors propose a DRL-based PPO framework to optimize EE in LoRa networks, utilizing a single GW for data collection to extend the network lifetime.
In this work, we focus on optimizing key resource allocation parameters in LoRa systems, including SF allocation, TP assignment, bandwidth (W) selection, and EDs association, to maximize the global system EE. Our approach incorporates air-to-ground (A2G) link characteristics to account for the impact of the device-to-gateway channel on link quality. To the best of the authors’ knowledge, this is the first work to propose a multi-agent reinforcement learning (MARL) approach that uses multi-agent proximal policy optimization (MAPPO) to enhance global system EE in flying LoRa GWs. The main contributions of this paper are summarized as follows:
@@ II System Model
II-A EDs Mobility
To accurately model the mobility behavior of our LoRa ground devices, we initially deployed the EDs at random positions within the defined target area $S$. For a given ED $v$, a unique speed vector is assigned $\mathbf{s}_{v}(t)=[s_{v}^{x}(t),s_{v}^{y}(t)]$, where $s_{v}^{x}(t)$ and $s_{v}^{y}(t)$ are respectively the speed along $x$ and $y$ axes. Furthermore, the random assignment ensures that each ED moves in an independent direction with a distinct speed. At each time step $t\in\mathcal{T}=\{1,\ldots,T\}$, the EDs update their positions based on their assigned speed. If an ED reaches the area’s boundary, it bounces back by reversing its speed direction while ensuring it remains within the limits. Additionally, there is a $p_{0}$ probability per ED per time step that its speed will be randomly reassigned, introducing variability in movement over time.
@@ II-B Air-to-Ground Channel Model
In UAV communication systems, A2G link propagation plays a crucial role and can be leveraged for flying GWs communication [10]. In this work, a probabilistic path loss model was used to model A2G communication links, whereby the line-of-sight (LoS) and non-line-of-sight (NLoS) links were considered separately with different probabilities of occurrence. Consequently, the likelihood of having a LoS connection between GW $u$, and an ED $v$ is given by:
@@
where $\hat{\phi}$ = $sin^{-1}(\frac{h_{u}}{d_{u,v}})$ is the angle of elevation from GWs $u$, $u\in\mathcal{U}$ to EDs $v$, $v\in\mathcal{V}$. $h_{u}$ is the altitude of the UAV $u$. The Euclidean distance between $u$ and $v$ is denoted as $d_{u,v}$. $\vartheta$ and $\lambda$ are constants that depend on the environment. Note, from Eq. (1), the probability values for $P_{NLoS}$ can be expressed as $P_{NLoS}(\hat{\phi})=1-P^{LoS}(\hat{\phi})$. Furthermore, the path loss model from UAV $u$ to the ground LoRa ED $v$ can be expressed as:
@@ II-C Energy Efficiency Model
where $W_{v}(t)$ is the bandwidth allocated to the communication link.
Furthermore, the EE of each UAV is calculated by dividing the sum of all uplink data rates from all EDs connected to the UAV by the total consumed power. Hence, we define the EE $\zeta_{u}(t)$ of UAV $u$ at time $t$ as:
@@ II-D EDs Association and Resource Allocation Scheme
In the UAV-mounted LoRa GWs system, implementing a dynamic EDs association is a crucial element in efficiently managing connections between ground EDs and flying GWs. Given that each UAV serves multiple EDs, it is important to ensure the association is properly established to ensure load balance in the network. Therefore, the binary association between a UAV $u$ and an ED $v$ at time slot $t$ is denoted by $a_{u,v}(t)\in\{0,1\}$, $a_{u,v}(t)=1$ if ED $v$ is being served by UAV $u$ at time $t$, and $a_{u,v}(t)=0$ otherwise. The index of the UAV selected by ED $v$ at time $t$ can be expressed as $a_{v}(t)=\sum_{u\in\mathcal{U}}a_{u,v}(t)\cdot u.$ [^1] This scalar value represents the UAV associated with ED $v$ and is essential for calculating interference in the SINR as given in Eq. (5).
@@
$\sum_{j=1}^{\mathcal{J}}p_{v,j}(t)\leq 1,\forall v\in\mathcal{V}.$ (9)
@@ III Problem Formulation
This section presents a detailed description of the optimization problem based on our proposed system model. The objective of the paper is to maximize the global system EE in flying LoRa GWs, as defined in Eq.(11a). Consequently, our optimization problem is formulated as follows.
@@
[…]
$\;a_{u,v}(t)\in\{0,1\},\forall u,v,$ (11c)
$\;\sum_{u\in\mathcal{U}}a_{u,v}(t)\leq 1,\forall v,$ (11d)
$\;\sum_{v\in\mathcal{V}}a_{u,v}(t)\leq\Lambda_{max},\forall u,$ (11e)
$\;\psi_{v}(t)\in{\bf\Psi},\forall v,$ (11f)
$\;\mho_{u,v}\geq\bar{\mho}_{SF},\forall u,v,$ (11g)
$\;w_{v,m}(t)\in\textbf{W},\forall v,$ (11h)
[…]
Constraint  (11b) limits the lower and the upper bounds of the transmission power available for each LoRa ED. The binary indicator of ED $a_{u,v}$ in  (11c), is either 0 or 1. In constraint  (11d), each ED can access at most one channel. Constraint  (11e) ensures that at most, there are $\Lambda_{max}$ EDs $v$ that can be associated to a single gateway $u$. Furthermore, constraint  (11f) specifies the available set of SFs. It is important to note that constraint  (11g) ensures that the received SINR must not be smaller than the threshold $\bar{\mho}$ shown in Table I, to detect the required LoRa EDs adopting a particular SF correctly [6] and finally, constraint  (11h) ensures that the bandwidth selected for each ED $v$ is from the discrete set W.
TABLE I: SINR Thresholds ¯$\bar{\mho}_{SF}$ with $W=125$ kHz [6]
SF 7 8 9 10 11 12
¯$\bar{\mho}_{SF}$ (dB) -7.5 -10 -12.5 -15 -17.5 -20
The formulated optimization problem (P1) is NP-hard due to its combination of binary and stochastic constraints, along with the selection of discrete variables for SF, TP, and W, leading to a non-convex objective function that is challenging for conventional optimization methods. As established in [12], no efficient polynomial-time solution exists for this problem. Consequently, traditional RL methods, which require storing all MDP tuples in a table, are computationally intensive and unsuitable for multi-agent scenarios; thus, we design POMDP method and leverage MAPPO-based algorithm to tackle the formulated problem in Eq.(11a).
@@ IV Multi-Agent PPO Algorithm
MAPPO is a promising framework that builds on the centralized training, decentralized execution (CTDE) concept and extends the PPO to the multi-agent system [13, 14]. It focuses on the actor-critic architecture, which consists of two components: the actor, who is in charge of making decisions, and the critic, who analyses those actions using a value function. In a CTDE architecture shown in Fig. 2, the critic network learns a centralized value function that includes knowledge of all agents’ activities, whilst each agent uses the actor-network to determine its policy based solely on local observations. Furthermore, due to the limited communication range and the maximum number of EDs each flying LoRa GW can associate with at any given time, the global state of the system cannot be fully observed by a single agent. Consequently, we model the problem using a partially observable Markov decision process (POMDP). In this framework, each agent only has partial information about the environment, leading to uncertainty in decision-making. Formally, the POMDP is defined by the tuple $\left\langle\mathcal{U},\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma,\mathcal{O}\right\rangle$.
@@
2) States $\mathcal{S}$: The global state set $\mathcal{S}(t)$ represents the environment’s current conditions at time step $t$. The global state is given by $\mathcal{S}(t)=\{\Psi,P,W,\Re,assEDs_{pos},GWs_{pos}\}$. These elements correspond to the assigned SF, TP, bandwidth, the computed data rate, the 2D positions of the associated EDs, and the positions of the GWs.
3) Observations $\mathcal{O}$: Due to partial observability, each flying LoRa GW $u$ at time step $t$ perceives only a subset of the global state $\mathcal{S}(t)$. Therefore, the observation space $o_{u}(t)\in\mathcal{S}$ for agent $u$ is defined as $o_{u}(t)=\{\Psi_{v}(t),P_{v}(t),\Re_{v}(t),assEDs_{pos,v}(t),GWs_{neighbors}(t)\}$. Here, the set $assEDs_{pos,v}$ represents the positions of the EDs currently associated with GW $u$. This partial observation mechanism ensures that each GW makes decisions based on local information, requiring cooperative strategies for optimal network-wide performance.
4) Actions $\mathcal{A}$: At each time step $t$, each agent makes decisions regarding resource allocation. These actions include selecting the appropriate SF $\psi(t)$, TP $p(t)$, and bandwidth $w(t)$. The actions taken by an agent can be expressed as $A_{u}(t)=(\psi(t),p(t),w(t))$. Once a decision has been made, the selected LoRa ED is served using the allocated resources.
5) Rewards $\mathcal{R}$: The reward function is related to the global EE system of all flying LoRa GWs, as described in 11a. There is defined as:
$r_{u}(t)=\sum_{t=1}^{T}\sum_{u=1}^{U}\frac{\sum_{v\in\mathcal{V}}\Re_{u,v}(t)a_{u,v}(t)}{\sum_{v\in\mathcal{V}}P_{u,v}(t)a_{u,v}(t)+P_{c}}.$ (12)
We ensure that the constraints in (11f) and (11g) are respected before rewards are assigned to the agents. Specifically, when the observation ${o}_{u}(t)$ is received, the agent must take an action ${a}_{t}\in\mathcal{A}$ that satisfies the constraints in P1.
5) Policy function: The policy function $\boldsymbol{\pi_{\theta_{u}}}(\textbf{a}_{u},o_{u})$ is modeled by an actor-network, with the vector $\boldsymbol{\theta}_{u}$ as its parameters. It determines the strategy for the flying GWs based on their local observations.
6) Value function: The value function $V_{\phi}(\boldsymbol{o_{u}})$ is implemented by a critic network parameterized by $\phi$. This network evaluates the expected future rewards for flying GWs given the current state $\boldsymbol{o}_{u}$. The critic aims to learn a value function that approximates the optimal future rewards, guiding the agents toward the globally optimal policy that maximizes their long-term rewards.
Figure 2: Illustration of Centralized Critic and Decentralized Actor for MAPPO training.

### v2

@@ Abstract
As next-generation Internet of Things (NG-IoT) networks continue to grow, the number of connected devices is rapidly increasing, along with their energy demands. This creates challenges for resource management and sustainability. Energy-efficient communication, particularly for power-limited IoT devices, is therefore a key research focus. In this paper, we deployed flying LoRa gateways mounted on unmanned aerial vehicles (UAVs) to collect data from LoRa end devices and transmit it to a central server. Our primary objective is to maximize the global system energy efficiency of wireless LoRa networks by joint optimization of transmission power, spreading factor, bandwidth, and user association. To solve this challenging problem, we model the problem as a partially observable Markov decision process (POMDP), where each flying LoRa GW acts as a learning agent using a cooperative multi-agent reinforcement learning (MARL). Simulation results demonstrate that our proposed method, based on the multi-agent proximal policy optimization (MAPPO) algorithm, significantly improves the global system EE and surpasses the conventional MARL schemes.
@@ I Introduction
The next-generation Internet of Things (NG-IoT) technologies for 5G and 6G applications are revolutionizing communication by enabling seamless data exchange between devices and networks. This evolution drives intelligent applications in areas such as healthcare, smart cities, agriculture, and autonomous vehicles [1]. With IoT device connections expected to reach 125 billion by 2030 [2], energy consumption presents a significant challenge, particularly in maintaining low-power and long-range communication networks. Enhancing energy efficiency (EE) is therefore essential to align with global sustainability objectives, including the United Nations Sustainable Development Goal 7: Affordable and Clean Energy [3].
Low-power wide area networks (LPWANs), particularly long-range (LoRa) technology, have emerged as a cost-effective solution for long-distance communication in IoT applications. However, existing terrestrial LoRa networks depend on fixed ground-based gateways (GWs), which struggle with non-line-of-sight (NLoS) propagation, especially in dense urban or remote environments. While deploying additional terrestrial LoRa GWs is affordable, it does not necessarily resolve NLoS issues. Conversely, satellite-based IoT solutions, such as the FOSSA system [^1] FOSSA systems is a low-Earth orbit (LEO) satellite network providing global IoT connectivity for remote areas. More details at: https://fossa.systems/, aim at connecting IoT devices to non-terrestrial networks. However, this approach significantly increases transmission power requirements and introduces higher latency, making it impractical for many energy-constrained IoT applications.
Beyond infrastructure deployment, effective resource allocation plays a crucial role in optimizing LoRa network performance. Existing studies primarily rely on alternative optimization techniques, where complex optimization problems are decomposed into sub-problems and solved iteratively [4]. Although alternative optimization methods have proven effective in certain network environments, they often struggle to adapt dynamically to varying IoT traffic patterns and environmental conditions, leading to suboptimal resource utilization. For example, the authors in [5] employ an alternative optimization-based method for a single-flying LoRa GW, which lacks adaptability in dynamic environments. In contrast, reinforcement learning (RL)-based approaches, such as those in [6, 7], utilize the Q-learning approach for resource allocation. However, these methods depend on static Q-tables, making them impractical for managing complex and dynamic IoT environments. To improve the EE system of LoRa, the work in [8] proposes a framework based on deep reinforcement learning (DRL) proximal policy optimization (PPO), but it is limited to a single GW for data collection, which restricts scalability and flexibility.
To address these limitations, we propose a UAV-assisted multi-flying LoRa GW deployment. This novel approach can dynamically reposition GWs to enhance network coverage, mitigate NLoS issues, optimize end-devices (EDs) association, minimize transmission power, and maximize energy-efficient communication. Furthermore, we utilize a multi-agent reinforcement learning (MARL) framework with multi-agent PPO (MAPPO) to optimize resource allocation in multi-flying LoRa GWs. Although single-agent RL approaches have shown promise in simpler scenarios [8], extending them to multi-UAV LoRa networks introduces significant challenges, including partial observability, decentralized control, and the need to optimize multiple interdependent parameters in dynamic environments. Our MAPPO approach overcomes these challenges by allowing UAV-mounted GWs to autonomously manage key network parameters, including spreading factor (SF) allocation, transmission power (TP) control, bandwidth (W) selection, and ED association. By incorporating air-to-ground (A2G) link characteristics, our framework ensures efficient resource distribution while maximizing system-wide EE. Unlike traditional methods, MAPPO continuously updates policies in real time, making it more adaptable to dynamic network conditions. To the best of the authors’ knowledge, this is the first work to propose a MARL-based approach for optimizing global system EE in a multi-flying LoRa GW deployment. The main contributions of this paper are summarized as follows:
@@ II System Model
II-A LoRa End Devices Mobility
To accurately model the mobility behavior of our LoRa ground EDs, we initially deployed the EDs at random positions within an area of interest. For a given ED $v$, a unique speed vector is assigned $\mathbf{s}_{v}(t)=[s_{v}^{x}(t),s_{v}^{y}(t)]$, where $s_{v}^{x}(t)$ and $s_{v}^{y}(t)$ are respectively the speed along $x$ and $y$ axes. Furthermore, the random assignment ensures that each ED moves in an independent direction with a distinct speed. At each time step $t\in\mathcal{T}=\{1,\ldots,T\}$, the EDs update their positions based on their assigned speed. If an ED reaches the area’s boundary, it bounces back by reversing its speed direction while ensuring it remains within the limits. Additionally, there is a $p_{0}$ probability per ED per time step that its speed will be randomly reassigned, introducing variability in movement over time.
Figure 1: The studied system model.
@@ II-B Air-to-Ground Channel Model
In UAV communication systems, A2G link propagation plays a crucial role and can be leveraged for flying GWs communication [9]. In this work, a probabilistic path loss model was used to model A2G communication links, whereby the line-of-sight (LoS) and NLoS links were considered separately with different probabilities of occurrence. Consequently, the likelihood of having a LoS connection between GW $u$, and an ED $v$ is given by:
@@
where $\hat{\phi}$ = $sin^{-1}(\frac{h_{u}}{d_{u,v}})$ is the angle of elevation from GWs $u$, $u\in\mathcal{U}$ to EDs $v$, $v\in\mathcal{V}$. $h_{u}$ is the altitude of the UAV $u$. The Euclidean distance between $u$ and $v$ is denoted as $d_{u,v}$. $\vartheta$ and $\lambda$ are constants that depend on the environment. Note, from Eq. (1), the probability values for $P_{NLoS}$ can be expressed as $P_{NLoS}(\hat{\phi})=1-P_{LoS}(\hat{\phi})$. Furthermore, the path loss model from UAV $u$ to the ground LoRa ED $v$ can be expressed as:
@@ II-C Energy Efficiency Model
where $W_{v}(t)$ is the bandwidth allocated to the communication link. Furthermore, the EE of each UAV is calculated by dividing the sum of all uplink data rates from all EDs connected to the UAV by the total consumed power. Hence, we define the EE $\zeta_{u}(t)$ of UAV $u$ at time $t$ as:
@@ II-D EDs Association and Resource Allocation Scheme
In the UAV-mounted LoRa GWs system, implementing a dynamic EDs association is a crucial element in efficiently managing connections between ground EDs and flying GWs. Given that each UAV serves multiple EDs, it is important to ensure the association is properly established to ensure load balance in the network. Therefore, the binary association between a UAV $u$ and an ED $v$ at time slot $t$ is denoted by $a_{u,v}(t)\in\{0,1\}$, $a_{u,v}(t)=1$ if ED $v$ is being served by UAV $u$ at time $t$, and $a_{u,v}(t)=0$ otherwise. The index of the UAV selected by ED $v$ at time $t$ can be expressed as $a_{v}(t)=\sum_{u\in\mathcal{U}}a_{u,v}(t)\cdot u$.
@@
$\sum_{j=1}^{\mathcal{J}}p_{v,j}(t)\leq 1,\forall v\in\mathcal{V},t\in\mathcal{T}.$ (9)
@@ III Problem Formulation
This section provides a detailed description of the optimization problem based on our proposed system model. The objective of this paper is to maximize the global system EE of flying LoRa GWs, as defined in Eq. (11a). Consequently, we formulate our optimization problem as follows:
$\max_{\mathbf{\bar{\Psi},\bar{P},a,\bar{W}}}$ $\quad\sum_{t=1}^{T}\sum_{u=1}^{U}\zeta_{u}(t),$ (11a)
s.t. $\;a_{u,v}(t)\in\{0,1\},\forall u,v,t\in\mathcal{T},$ (11b)
$\;\sum_{u\in\mathcal{U}}a_{u,v}(t)\leq 1,\forall v,t\in\mathcal{T},$ (11c)
$\;\sum_{v\in\mathcal{V}}a_{u,v}(t)\leq\Lambda_{max},\forall u,t\in\mathcal{T},$ (11d)
$\;\psi_{v,n}(t)\in{\bf\Psi},\forall v,t\in\mathcal{T},$ (11e)
$\;p_{v,j}\in\textbf{P},\forall v,$ (11f)
$\;w_{v,m}(t)\in\textbf{W},\forall v,t\in\mathcal{T},$ (11g)
$\;\rho_{u,v}(t)\geq\text{SNR}_{threshold},\forall u,v,t\in\mathcal{T}$ (11h)
@@
Here, the constraint  (11b) defines the binary indicator $a_{u,v}$, which specifies whether an ED is associated with a GW. Constraint  (11c) ensures that each ED can access at most one channel. Constraint  (11d) limits the number of EDs $v$ that can be associated with a single GW $u$ to at most $\Lambda_{max}$. Furthermore, constraint  (11e) defines the available set of SFs. Constraint  (11f) restricts the transmit power of the ED to be selected from a predefined discrete set. Constraint  (11g) ensures that the bandwidth allocated to each ED $v$ is chosen from a discrete set W. Finally, constraint  (11h) enforces that the received SNR must not fall below the threshold $\text{SNR}_{threshold}$, as shown in Table I, to ensure correct detection of LoRa EDs adopting a specific SF [6].
The formulated optimization problem (11) is NP-hard due to its combination of binary and stochastic constraints, along with the selection of discrete variables for SF, TP, and W, leading to a non-convex objective function that is challenging for conventional optimization methods. As established in [11], no efficient polynomial-time solution exists for this problem. Consequently, traditional RL methods, which require storing all MDP tuples in a table, are computationally intensive and unsuitable for multi-agent scenarios; thus, we design POMDP method and leverage the MAPPO-based algorithm to tackle the formulated problem in Eq.(11).
@@ IV Multi-Agent PPO Algorithm
MAPPO is a promising framework that builds on the centralized training, decentralized execution (CTDE) scheme and extends the PPO to the multi-agent system [12, 13]. It focuses on the actor-critic architecture, which consists of two components: the actor, who is in charge of making decisions, and the critic, who analyses those actions using a value function. In a CTDE architecture shown in Fig. 2, the critic network learns a centralized value function that includes knowledge of all agents’ activities, whilst each agent uses the actor-network to determine its policy based solely on local observations. Furthermore, due to the limited communication range and the maximum number of EDs each flying LoRa GW can associate with at any given time, the global state of the system cannot be fully observed by a single agent. Consequently, we model the problem using a POMDP. In this framework, each agent only has partial information about the environment, leading to uncertainty in decision-making. Formally, the POMDP is defined by the tuple $\left\langle\mathcal{U},\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma,\mathcal{O}\right\rangle$.
@@
2) States $\mathcal{S}$: The global state represents the complete environment configuration at time step $t$. Therefore, we define the global state vector of all UAVs as $\mathbf{\mathcal{S}}(t)=\left\{s_{1}(t),s_{2}(t),\dots,s_{u}(t),\dots,s_{U}(t)\right\}$, where $s_{u}(t)$ denotes the local state of agent $u\in\mathcal{U}$ at time $t$.
3) Observations $\mathcal{O}$: Due to partial observability, each flying LoRa GW $u$ at time $t$ perceives only a subset of the global state $\mathbf{\mathcal{S}}(t)$. Hence, the local observation $o_{u}(t)\in\mathcal{O}_{u}\subset\mathcal{S}(t)$ for agent $u$ is denoted as $o_{u}(t)=\left\{\psi(t),p(t),\Re(t),assEDs_{pos}(t),GWs_{pos}(t)\right\}$. Here, $assEDs_{pos}(t)$ represents the positions of the EDs currently associated with GW $u$, and $GWs_{pos}(t)$ are the positions of neighboring GWs within communication range. This partial observation mechanism ensures that each GW makes decisions based on its local view of the environment.
TABLE I: SINR Thresholds with $W=125$ kHz [6]
SF 7 8 9 10 11 12
SNRthreshold (dB) -7.5 -10 -12.5 -15 -17.5 -20
4) Actions $\mathcal{A}$: Each agent $u\in\mathcal{U}$ selects an action $a_{u}(t)\in\mathcal{A}_{u}$ at time $t$. The action comprises the selection of communication parameters including spreading factor, transmission power, and bandwidth. This can be represented as $a_{u}(t)=\left\{\psi_{u}(t),p_{u}(t),w_{u}(t)\right\}$, where $\psi_{u}(t)$, $p_{u}(t)$, and $w_{u}(t)$ denote the SF, TP, and bandwidth chosen by agent $u$ at time $t$, respectively.
5) Reward Function $\mathcal{R}$: The reward for each flying LoRa GW $u$ at time step $t$ is defined as:
$r_{u}(t)=\frac{\sum\limits_{v\in\mathcal{V}}\Re_{u,v}(t)\cdot a_{u,v}(t)}{\sum\limits_{v\in\mathcal{V}}P_{u,v}(t)\cdot a_{u,v}(t)+P_{c}}.$ (12)
Note that the reward in Eq.(12) is allocated only if the SNR constraint (11h) is satisfied; otherwise, the reward is zero. This constraint can be formulated as:
[…]
Therefore, the cumulative reward over the entire time horizon $T$ and all GWs $\mathcal{U}$ is:
$\mathcal{R}=\sum_{t=1}^{T}\sum_{u=1}^{U}r_{u}(t)$ (14)
6) Policy function: The policy function $\boldsymbol{\pi_{\theta_{u}}}(\textbf{a}_{u},o_{u})$ is modeled by an actor-network, with the vector $\boldsymbol{\theta}_{u}$ as its parameters. It determines the strategy for the flying GWs based on their local observations.
7) Value function: The value function $V_{\phi}(\boldsymbol{o_{u}})$ is implemented by a critic network parameterized by $\phi$. This network evaluates the expected future rewards for flying GWs given the current state $\boldsymbol{o}_{u}$. The critic aims to learn a value function that approximates the optimal future rewards, guiding the agents toward the globally optimal policy that maximizes their long-term rewards.